# IntentRec: Predicting User Session Intent with Hierarchical Multi-Task Learning

### v1

@@ 1. Introduction
The exact definition of a user’s intent varies across diverse applications and is often hidden. In Netflix, we identify several interaction metadata that can be associated with the user intent. Specifically, we leverage the type of action a member takes on the product as a direct reflection of what they intend to do on the platform. For example, when a member plays a follow-up episode or scene of something they were already watching previously can be categorized as “continue watching” intent. Additionally, intents can also be whether a member wants to watch a movie or TV show, or is in the mood for a short-watch session or a long binge-watching session. As shown in Fig. 1, we have different metadata of user interactions that can be mapped to intents including, Action Type (e.g., discovering new content vs binge watching, etc.), Genre (e.g., horror, thriller, drama, etc.), movie or TV show preference, etc. While predicting the next item ID is the most important task, anticipating the user’s future intent (e.g., action type prediction) is also crucial, as it can enhance the next-item prediction.
@@ 3.2. Input Feature Sequence Formation
We assume user engagement data consists of historical interactions (e.g. clicks, plays, etc.) that users have made in Netflix. We use the latest $n$ (e.g., 100) interactions of each user for training and testing. A user $u$’s engagement is represented as a temporal interaction sequence $\{int_{1},\ldots,int_{n}\}$ (latest at the end).
Each interaction $int_{k},1\leq k\leq n$ is associated with various metadata. For instance, categorical features include item-ID $i_{k}$, action type $c_{k}$ (see Section 4.1.1 for details), genre $g_{k}$, and numerical features include timestamp $t_{k}$, episode position $e_{k}$, etc. We convert all categorical metadata of an interaction $int_{k}$ into embedding using a trainable embedding layer per metadata. Among all categorical features, item-ID has the largest embedding dimension (e.g., 400), and the others have relatively small dimensions (e.g., 20). Regarding numerical metadata, they are normalized first (e.g., between 0 and 1) and become 1-dimensional features. Some interaction metadata are used as both intent prediction labels and input features. Different interaction metadata would be used in other datasets/applications.
@@ 3.3. User Intent Prediction
Figure 4. Given an input feature sequence of a user, a user intent embedding sequence is constructed by an attention-based aggregation of auxiliary prediction (e.g., Action Type and Genre) results.
@@ 4.1.1. Dataset
Table 1. Summary of Netflix user engagement dataset.
Number of training/
validation/test users
Number of
Items
Interaction
Metadata
2.2M / 181K / 176K 35K
Action Type, Genre,
Movie/Show Tag,
Timestamps, Duration, $\cdots$
Action Type
Labels
Genre
Labels
Movie/Show
Labels
Time-since-
release Labels
11 (e.g.,
Rewatch $\rightarrow$ 1)
21 (e.g.,
Comedy $\rightarrow$ 1)
2 (e.g.,
Movie $\rightarrow$ 1)
3 (e.g.,
$\leq$ 1 week $\rightarrow$ 1)
Table 2. Next-item and next-intent prediction results of baselines and our proposed method IntentRec on the Netflix user engagement dataset. All the metrics are represented as relative % improvements compared to the LSTM (Hochreiter and Schmidhuber, 1997) baseline. N/A indicates that a model is not capable of predicting a certain intent (e.g., Action Type) of a user. The best baseline results are colored gray. For the next-item prediction task, IntentRec presents 7.4% and 5.6% improvements in MRR and WMRR compared to the best baseline, with statistical significance.
@@
LSTM (Hochreiter and Schmidhuber, 1997) (+0.00%, +0.00%) +0.00% +0.00% +0.00% +0.00%
GRU (Cho et al., 2014) (+4.45%, +4.30%) +0.12% +0.18% +0.01% +0.07%
Transformer (Vaswani et al., 2017) (+10.3%, +8.41%) +2.35% +0.94% +0.16% +0.17%
SASRec (Kang and McAuley, 2018) (+6.62%, +6.20%) N/A N/A N/A N/A
BERT4Rec (Sun et al., 2019) (+0.25%, -1.97%) N/A N/A N/A N/A
TransAct (Xia et al., 2023) (+33.7%, +30.6%) +8.12% +2.54% +0.94% +0.58%
@@
(+43.6%, +37.9%) +11.7% +5.39% +1.35% +1.42%
We use proprietary user engagement data collected in Netflix. We preprocess raw engagement sequences (i.e., every interactions a user has made on Netflix) of users. The preprocessing is based on some assumptions gathered from internal research to create higher-level engagements. We then randomly sample users from all users to generate training/validation/test data (split ratio: 86%/7%/7%). Key statistics of the Netflix dataset is given in Table 1.
Rich metadata of each interaction is available in the Netflix dataset; e.g., ‘Action Type’ information indicates a category of an interaction such as continuing to watch something the user has previously started watching, discovering content that the member has never seen on Netflix before, and others. Among all metadata, we select 4 key metadata (Action type, Genre preference, Movie/Show type preference, and Time-since-release) to include as intent prediction labels, as (1) they are considered to be the most important proxies that can help determine a member’s latent intent, and (2) predicting them has high business values in several downstream applications. Note that these intent prediction labels are subject to change depending on datasets/applications. Detailed information about each intent prediction label is given as follows.
  1. (1)
    Action Type: 11 unique categories of user interactions made on Netflix such as discovering new content, binge-watching existing shows, etc. These labels are created by a rule-based algorithm. We define ”core” action types as 5 distinct types (from New Content Watch to Rewatch) related to actual playing and focus on predicting core labels.
  2. (2)
    Genre: 21 pre-defined genre labels such as comedy, action, etc. Users’ genre preferences are diverse, and predicting genres can help us capture users’ hidden intents precisely.
  3. (3)
    Movie/Show: a binary label indicating the corresponding item is a movie or a show. Predicting it is crucial as users may have already decided whether they want to watch shows (relatively short duration) or movies.
  4. (4)
    Time-since-release: 3 manually-chosen labels indicating whether the corresponding item is released recently or not; 1 = within a week, 2 = between a week and a month, and 3 = older than a month. Predicting it is related to discovering users’ temporal preferences (e.g. a user enjoying the newly released shows vs old-fashioned films).
Following the experimental setup used in existing literature (Pal et al., 2020; Xia et al., 2023), all our experiments are based on the Netflix dataset. We do not use public datasets in our experiments, since they lack of crucial interaction metadata (e.g., action type, interaction duration, etc.), which are needed for intent and item predictions in IntentRec.
@@ 4.1.2. Baselines
We use the following state-of-the-art sequential recommendation models as baselines. We have excluded baselines (Chen et al., 2019) having similar or older architectures than our current baselines or that are not sequential recommenders (Zhou et al., 2019). Note that all the baselines except TransAct (Xia et al., 2023) and IntentRec-V0 are not capable of employing interaction metadata as features for training and testing.
@@ 4.1.3. Evaluation Metrics
We use various standard ranking metrics typically used in recommender systems for evaluation but for brevity, we report mean reciprocal rank (MRR) (Craswell, 2009) and weighted MRR (WMRR). We focus on these two metrics since other metrics (e.g., Recall, NDCG, AUC) show similar tendencies. The MRR and WMRR metrics indicate how accurately a model predicts the next item $i_{n+1}$ of a test user $u$, given her $n$ observed items $\{i_{1},\ldots,i_{n}\}$. MRR and WMRR metrics are defined as follows.
(5) $MRR=\frac{1}{|Test|}\sum_{u\in Test}\frac{1}{R(i_{n+1})},WMRR=\frac{\sum_{u\in Test}\frac{d_{n+1}}{R(i_{n+1})}}{\sum_{u\in Test}d_{n+1}},$
where $d_{i}$ is the duration of an interaction (e.g., playtime), and $R(i)$ is a predicted rank of an item $i$ among all items. For all experiments, we aim to predict the latest item of a user; hence, $n$ is set to the maximum sequence length (100). Due to the proprietary nature of the paper, we only provide relative % improvements of MRR and WMRR of a method compared to the worst baseline (e.g., LSTM) instead of the absolute values, while presenting experimental results.
@@
Table 4. Ablation study of each prediction head of IntentRec on the Netflix user engagement dataset. All the metrics are represented as relative % improvements compared to the V0 baseline for item-ID prediction and each variant for intent prediction. N/A indicates that a model is not capable of predicting a certain intent (e.g., Action Type) of a user. Action Type prediction is the most important one for next-item prediction.
@@
4.1.4. Implementation Details and Hyperparameters
IntentRec is developed in Python and Tensorflow with Keras and tested in the Amazon EC2 P3 machine with 8 Testla-V100 GPUs (16GB memory). We cannot release the code and dataset used in IntentRec due to the proprietary nature of those artifacts. Hyperparameters of IntentRec are tuned by grid searches on the validation data; specifically, the maximum sequence length $n$ is set to 100, the dimensions of interaction feature, short-term embedding, and intent representation are set to $523$, $200$, $200$, respectively, and the output dimensions of the Transformer intent and item encoder are set to $600$ and $800$. Each Transformer consists of two encoder layers equipped with a $512$-dimensional fully connected layer and $8$ attention heads. We also use timestamp-based positional encoding for the Transformer encoder. We train IntentRec with Adam (Kingma and Ba, 2015) optimizer for $10$ epochs with a learning rate: $0.0005$ and batch size: $1024$. The intent prediction coefficient $\lambda$ is set to $1.0$.
@@ 4.2. Next Item and Intent Prediction Accuracy
As shown in Table 2, IntentRec presents the highest next-item prediction accuracy across all baselines, as per both MRR and WMRR metrics. Remarkably, IntentRec outperforms the best baselines: TransAct and IntentRec-V0; for instance, IntentRec shows 7.4% and 5.6% improvements in MRR and WMRR compared to TransAct with statistical significance (p-values from Student’s t-test $<0.01$). Regarding baselines, they all exhibit limited next-item prediction performance as they cannot predict users’ intents (marked as N/A in the table) or incorporate the user intents to the next-item prediction directly. TransAct (Xia et al., 2023) and IntentRec-V0 show relatively higher accuracy than other baselines as they can leverage interaction metadata, while other baselines do not incorporate interaction metadata due to their architectural limitations.
IntentRec also shows superior prediction performance of user intents among all methods; IntentRec shows the highest MRR values across all intent prediction tasks (e.g., Action Type). However, the intent prediction accuracy improvements are smaller than the next-item prediction ones as the hierarchical learning of IntentRec prioritizes to optimize next-item prediction over next-intent predictions; in other words, the intent predictor of IntentRec can be fine-tuned to find the optimal solutions for next-item prediction, not for next-intent prediction. This trade-off can be moderated by changing the intent prediction coefficient $\lambda$ during the training (i.e., larger $\lambda$ focuses more on intent prediction). Refer to Fig. 8 for the impact of different $\lambda$ values on item and intent predictions.
@@ 4.4. Qualitative Analysis: User Intent Clustering
Given a user and her previous interactions, IntentRec is able to predict this user’s current intent and generate an intent embedding by aggregating auxiliary prediction results (e.g., Action Type, Genre) via an attention layer. We conduct a qualitative analysis of user intent embeddings to validate their quality and accuracy. Specifically, we apply K-means++ (Arthur and Vassilvitskii, 2007) clustering algorithm on intent embeddings of all users in the training data where the number of clusters (K) is set to 10. For each found cluster, we randomly sample a few users (e.g., 10) close to the cluster center on the embedding space. After that, we manually investigate the similarities between chosen users and determine the concept of the intent cluster (e.g., binge-watchers) by the commonalities between users.
@@ 5. Discussion
Computational overhead of Hierarchical MTL. The main computational bottleneck of IntentRec is the item predictor, since the total number of items (e.g., 35K) for predictions is much larger than the number of prediction labels for each intent (e.g., 2-20). Thus, the computational overhead from the input feature sequence formation and intent prediction module is marginal, compared to the next-item prediction cost. Investigating the optimization techniques for fast and scalable training is our future work.
@@
Hyperparameter Sensitivity. The key hyperparameters of IntentRec include the intent prediction strength $\lambda$, embedding dimensions of the intent prediction and short-term interest feature, and the output dimension of the Transformer item encoder. We test how much the prediction accuracy of IntentRec changes if we vary the values of key hyperparameters.
Fig. 8 shows hyperparameter sensitivity results of IntentRec on validation data, as per item and genre prediction MRR metrics. We omit the other intent prediction metrics such as Action Type or Movie/Show since they remain similar accuracy across different hyperparameters. Strong emphasis on intent prediction (e.g., $\lambda=1.0$) leads to both high item and intent prediction accuracy, as accurate intent prediction would act as a crucial hint for the item prediction. Regarding the embedding dimensions of the intent prediction and short-term interest feature, both intent and item accuracy do not change significantly across different values. As a large embedding dimension can slow down the training process, a medium value (e.g., 200) are recommended. Finally, the output dimension of the Transformer item encoder can directly affect the final performance of IntentRec, as the item prediction is our primary task. If the main focus is faster training and inference, a small output dimension (e.g., 400) is recommended. A large output dimension (e.g., 800) is needed to achieve the highest accuracy of IntentRec or obtain precise model parameters for downstream applications. We also conducted studies about embedding dimensions of different metadata, and in fact, embedding dimensions do not have a significant impact on the next-item prediction performance. Surprisingly, some of our most impactful features are just scalars (e.g., timestamps).
Figure 8. Hyperparameter sensitivity of IntentRec.

### v2

@@ 1. Introduction
The exact definition of a user’s intent varies across diverse applications and is often hidden. In Netflix, we identify several interaction metadata that can be associated with the user intent. Specifically, we leverage the type of action a member takes on the product as a direct reflection of what they intend to do on the platform. For example, when a member plays a follow-up episode or scene of something they were already watching previously can be categorized as “continue watching” intent. Additionally, intents can also be whether a member wants to watch a movie or TV show, or is in the mood for a short-watch session or a long continue-watching session. As shown in Fig. 1, we have different metadata of user interactions that can be mapped to intents including, Action Type (e.g., discovering new content vs continue watching, etc.), Genre (e.g., horror, thriller, drama, etc.), movie or TV show preference, etc. While predicting the next item ID is the most important task, anticipating the user’s future intent (e.g., action type prediction) is also crucial, as it can enhance the next-item prediction.
@@ 3.2. Input Feature Sequence Formation
We assume user engagement data consists of historical interactions (e.g. clicks, plays, etc.) that users have made in Netflix. We use the latest $n$ interactions of each user for training and testing. A user $u$’s engagement is represented as a temporal interaction sequence $\{int_{1},\ldots,int_{n}\}$ (latest at the end).
Each interaction $int_{k},1\leq k\leq n$ is associated with various metadata. We convert all categorical metadata of an interaction $int_{k}$ into embedding using embedding layers. Regarding numerical metadata, they are normalized first (e.g., between 0 and 1) and become 1-dimensional features. Some interaction metadata are used as both intent prediction labels and input features. Different interaction metadata would be used in other datasets/applications.
@@ 3.3. User Intent Prediction
Figure 4. Given an input feature sequence of a user, a user intent embedding sequence is constructed by an attention-based aggregation of auxiliary prediction (e.g., Action Type and Genre) results. We use ground-truth intent and item-ID labels to optimize predictions.
@@ 4.1.1. Dataset
Table 1. Next-item and next-intent prediction results of baselines and our proposed method IntentRec on the Netflix user engagement dataset. All the metrics are represented as relative % improvements compared to the TransAct (Xia et al., 2023) baseline. N/A indicates that a model is not capable of predicting a certain intent (e.g., Action Type) of a user. The best baseline results are colored gray. For the next-item prediction task, IntentRec presents 7.4% accuracy improvement compared to the best baseline, with statistical significance.
@@
LSTM -25.2% -7.51% -2.48% -0.93% -0.58%
GRU -21.9% -7.40% -2.30% -0.92% -0.51%
Transformer -17.5% -5.33% -1.56% -0.77% -0.41%
SASRec -20.3% N/A N/A N/A N/A
BERT4Rec -25.0% N/A N/A N/A N/A
TransAct +0.00% +0.00% +0.00% +0.00% +0.00%
@@
+7.40% +3.31% +2.78% +0.41% +0.84%
We use proprietary user engagement data collected in Netflix. We preprocess raw engagement sequences (i.e., every interactions a user has made on Netflix) of users. The preprocessing is based on some assumptions gathered from internal research to create higher-level engagements. We then randomly sample users from all users to generate training/validation/test data. Rich metadata of each interaction is available in the Netflix dataset; e.g., ‘Action Type’ information indicates a category of an interaction such as continuing to watch something the user has previously started watching, discovering content that the member has never seen on Netflix before, and others. Among all metadata, we select 4 key metadata (Action type, Genre preference, Movie/Show type preference, and Time-since-release) to include as intent prediction labels. Note that these intent prediction labels are subject to change depending on datasets/applications.
@@ 4.1.2. Baselines
We use the following state-of-the-art sequential recommendation models as baselines. We have excluded baselines (Chen et al., 2019) having similar or older architectures than our current baselines or that are not sequential recommenders (Zhou et al., 2019). Note that we added additional fully-connected layers to LSTM (Hochreiter and Schmidhuber, 1997), GRU (Cho et al., 2014), Transformer (Vaswani et al., 2017) baselines in order to predict user intent, while we used original implementations for other baselines.
@@ 4.1.3. Evaluation Metrics
We use various standard ranking metrics typically used in recommender systems for evaluation; for brevity, our “accuracy” metric is one of the standard ranking metrics (e.g., MRR, Recall, NDCG, AUC). Due to the proprietary nature of the paper, we only provide relative % improvements of accuracy of a method compared to the state-of-the-art baseline: TransAct (Xia et al., 2023).
@@
Table 3. Ablation study of each prediction head of IntentRec on the Netflix user engagement dataset. All the metrics are represented as relative % improvements compared to the V0 baseline for item-ID prediction and each variant for intent prediction. N/A indicates that a model is not capable of predicting a certain intent of a user. Action Type prediction is the most important one for next-item prediction.
@@ 4.2. Next Item and Intent Prediction Accuracy
As shown in Table 1, IntentRec presents the highest next-item prediction accuracy across all baselines. Remarkably, IntentRec outperforms the best baselines: TransAct and IntentRec-V0; for instance, IntentRec shows 7.4% accuracy improvement compared to TransAct with statistical significance (p-values from Student’s t-test $<0.01$). Regarding baselines, they all exhibit limited next-item prediction performance as they cannot predict users’ intents (marked as N/A in the table) or incorporate the user intents to the next-item prediction directly. TransAct (Xia et al., 2023) and IntentRec-V0 show relatively higher accuracy than other baselines as they can leverage interaction metadata, while other baselines do not incorporate interaction metadata due to their architectural limitations.
IntentRec also shows superior prediction performance of user intents among all methods; IntentRec shows the highest accuracy across all intent prediction tasks (e.g., Action Type). However, the intent prediction accuracy improvements are smaller than the next-item prediction ones as the hierarchical learning of IntentRec prioritizes to optimize next-item prediction over next-intent predictions; in other words, the intent predictor of IntentRec can be fine-tuned to find the optimal solutions for next-item prediction, not for next-intent prediction. This trade-off can be moderated by changing the intent prediction coefficient $\lambda$ during the training (i.e., larger $\lambda$ focuses more on intent prediction).
@@ 4.4. Qualitative Analysis: User Intent Clustering
Given a user and her previous interactions, IntentRec is able to predict this user’s current intent and generate an intent embedding by aggregating auxiliary prediction results (e.g., Action Type, Genre) via an attention layer. We conduct a qualitative analysis of user intent embeddings to validate their quality and accuracy. Specifically, we apply K-means++ (Arthur and Vassilvitskii, 2007) clustering algorithm on intent embeddings of all users in the training data where the number of clusters (K) is set to 10. For each found cluster, we randomly sample a few users (e.g., 10) close to the cluster center on the embedding space. After that, we manually investigate the similarities between chosen users and determine the concept of the intent cluster (e.g., Rewatchers) by the commonalities between users.